# Setting up for failure: automatic discovery of the neural mechanisms of cognitive errors

### v1

@@ 1 Introduction
Previous approaches have fallen short of these desiderata. For example, cognitive models – including drift-diffusion [Resulaj2009, PardoVazquez2019], Bayesian [Heald2021], symbolic [Castro2025], and large language model-based models [Binz2024] – have been successfully fit to detailed behaviour with impressive predictive power, but their architectures remained too abstract and divorced from RNNs to be useful as hypotheses about neural network dynamics (thus failing Requirement 1). Neural population coding models have suggested causal connections between neural response properties and particular patterns in behaviour [Matthey2015, Schneegans2017, McMaster2022], but remained mute about the neural network dynamics that give rise to the hypothesised neural responses in the first place (failing Requirement 1), have been rarely fit to behaviour quantitatively (failing Requirement 2), and their construction was not automated (failing Requirement 3). RNN models – either with hand-crafted [Edin2009, Bouchacourt2019], or with task-optimized architectures and parameters [Mante2013, Stroud2023] – have been highly successful at suggesting testable hypotheses about neural dynamics underlying cognitive performance, but they rarely capture behaviour beyond generically competent performance, let alone being quantitatively fit to detailed behavioural data (failing Requirements 2, and for hand-crafted networks also Requirement 3). In some cases, non-trivial aspects of behaviour were successfully captured by such models, but at the expense of including purpose-built design choices or ablations [Xie2023], thus making these models essentially hand-crafted in this sense (failing Requirement 3).
@@
In this work, we introduce an approach that satisfies all three requirements. We extend automated neural mechanism discovery by directly training biologically plausible RNNs to exhibit behavioural suboptimalities seen in experimental data. We tackle two major challenges for doing so. First, RNNs are data hungry, and behavioural data is scarce. To make training feasible, we generate synthetic data using a descriptive generative model, and use this to train our RNN (Figure 1). The second challenge is defining a suitable training objective for the RNN so that they can generate complex (for example, multimodal [Bays2009] or skewed [Fritsche2020, Resulaj2009]) continuous response distributions often found in experiments. Previously used methods for training RNNs are not appropriate for this. Task optimization of RNNs typically used simple optimisation criteria, such as mean squared error or cross entropy [Yang2019], which encourage unimodal, typically normal response distributions. Even when training RNNs explicitly to generate specific distributions of responses, moment- or score matching-based criteria were used [Echeveste2020, Chen2023], which also do not scale well to capturing complex distributions (and require arbitrary choices as to which moments or statistics of the target distribution are important). Instead, by framing elicitation of a behavioural output as generating samples in Euclidean space, we train the RNN using a training criterion inspired by diffusion models [Ho2020], a state-of-the-art class of generative neural models for complex, continuous distributions. We adapt the training procedure used for diffusion models to work within an RNN, harnessing their flexibility in sampling from complex distributions. This enables us to flexibly fit to continuous response data, unlike prior methods of fitting directly to behaviour [JiAn2025], which are limited to categorical distributions (§5).
@@
We demonstrate the power of our approach for automatic mechanism discovery on one of the most studied cognitive paradigms whose neural underpinning is still poorly understood: visual working memory (VWM). While there are many neural circuit models of working memory, they typically do not address the more challenging (and ecologically more relevant) situation when multiple pieces of information (‘items’) need to be maintained in memory Zhang1996. Even RNNs that do perform multi-item VWM tasks fail to capture the characteristic patterns of behaviour that are specific to the multi-item situation [Yang2019, Driscoll2022], or only capture some select behavioural biases by using hand-crafted networks with purpose-built architectural motifs with limited biological plausibility [Bouchacourt2019]. In particular, so-called swap errors are a major source of errors that arise when participants recall the wrong item from their working memory, instead of the item that has been cued – and thereby create strongly multimodal response distributions [Bays2009]. Current RNN models do not capture swap errors, and the resulting multimodal response distributions, at all. Conversely, there exist population coding models that do capture swap errors, but do not make predictions about neural circuit mechanisms; Schneegans2017). We address this gap by showing that our approach produces neural circuits that generate rich, realistic response distributions by directly training networks to perform swap errors. We use our trained networks to generate hypotheses about the biological implementation of observer models from the psychophysics literature (Schneegans2016, McMaster2022; §4), previously inaccessible to RNNs trained for task performance.
@@ 2 Background
where $\bm{r}$ is the vector of neural firing rates (hidden RNN activations), $\bm{s}$ is the input from an external population (e.g. sensory input into the population), $\bm{\eta}$ is the population’s biological process noise, and $\Lambda$ is the diagonal matrix of neurons’ membrane time constants (typically constant across cells, i.e. $\Lambda=\lambda I$). The non-linear computation $\mathcal{F}$, parameterised by $\theta_{r}$, models the recurrent communication between and information integration within neurons. The network output is generically $\bm{x}=\mathcal{G}(\bm{r};\theta_{o})$, often interpreted as a behavioural response. The exact form of $\mathcal{F}$ and $\mathcal{G}$ vary across studies; in this work, we use a dendritic tree architecture for each neuron (Appendix B; Lyo2024), and interpret the output of the network - a simple linear projection of the firing rates - as a colour estimate on the colour circle. More details are provided in §3, and discussion of our choice of architecture is provided in Appendix B - namely, we chose this architecture to be sufficiently expressive for the task posed, while also maintaining a biologically plausible connectivity structure.
Previous attempts to train RNNs that generate samples typically train RNNs on task-optimality or moment-matching criteria [Echeveste2020]. This is insufficient for us - as outlined in §1, we are interested in reproducing behaviour that takes on a distinctly multimodal distribution. To do so, we employ methods from the diffusion model literature. We briefly introduce diffusion models here, but defer many details to previous foundational work [Ho2020].
@@
Neural networks are data-hungry. This is not an obstacle training for task-optimality – task variables such as stimulus features and target responses can be generated procedurally – but it is when training directly on behaviour – producing human or animal behavioural datasets is prohibitively costly at the scales required for training networks. Therefore, we generate synthetic behavioural data to plug this gap. Many of the models listed in §1 which accurately describe behaviour using neural population codes [Schneegans2017], Bayesian cognitive [Heald2021] and descriptive [Bruijns2023] models, and/or large language models [Binz2024] can, by the same token, be used to faithfully generate synthetic behavioural data. These synthetic datasets can be used to train our dynamical neural model; sufficient flexibility in the generative model allows synthetic data that captures desired statistical facets and suboptimalities of behaviour. Many of these models, or their relatives, have also been used to fit behavioural data from working memory (WM) tasks, with varying degrees of satisfaction of our three requirements. For this reason, and because of the sophisticated response distributions elicited from even simple WM tasks, we choose this as the testbed for our method. In §5, we will return to potential extensions to other aspects of WM, and other cognitive tasks in general.
@@
While not consistently distinguished from short-term memory [Aben2012], WM is generally defined as the temporary maintenance, manipulation, and retrieval of sensory information over an order of seconds, for executive functions. A dominant experimental paradigm in investigating visual working memory (VWM) is the delayed estimation task, in which a human or animal subject is presented with sensory information and asked, after a short delay during which the stimulus is withdrawn, to reproduce some information about it. Figure 2 shows an example of a multiitem delay estimation task, where the subject is presented with multiple (here, 2) items, then presented with a probe dimension (here, location) feature value, called a cue, and asked to reproduce the corresponding report dimension (here, colour) feature value of the cued item. From here, we will interchangeably use ‘location’ and ‘probe feature’, and ‘colour’ and ‘report feature’ for ease of language, however the features used for each role is open for experimental design. Most responses fall near the cued item’s report feature value, and the distribution of estimation errors along the report feature (here, the colour circle) provides key evidence for measuring the capacity of VWM [Ma2014].
@@
However, there is also a central tendency in responses towards the colour of uncued items, a.k.a. distractors [Bays2009, Gorgoraptis2011]. This indicates the presence of swap errors, in which subjects mistakenly recall the colour of a distractor. Behavioural and neural evidence largely implicates misselection of the correct item from memory at cueing time, rather than a misbinding of feature values during memory encoding or storage – swap errors are a failure to extract the correct item from memory, rather than a misencoding into memory (but see Radmard2025) Macaque neural signatures reflects this misselection hypothesis Alleman2024. The key evidence for this is that swap errors are more likely to happen when a distractor’s probe feature value is close in feature space (here, physical space) to the cued one [Emrich2012, Schneegans2017, McMaster2022]. Describing the expected distribution of responses, including swap errors, requires a multimodal distribution, with modes placed over both target and uncued colours.
@@
We use BNS [Radmard2025] to generate synthetic multiitem delayed estimation data. This model predicts the likelihood of a swap error as a function of the distance of the distractor from the cued stimulus in each feature dimension (Appendix A). In the present work, we do not directly fit BNS to a real behavioural dataset. Instead, we opt to hand-tune the generative model to capture archetypal dependencies of swap errors. Specifically, we train RNNs with synthetic data generated by instantiations of BNS that either i) predicts no swap errors, ii) predicts the same frequency of swap errors for any distractor, and iii) predicts swap errors more frequently when its probe feature is more similar to that of the cued item, as per the extensive experimental evidence. In §3, we show that RNN representations resemble the real macaque neural geometry only in the third case, when maximal realism is captured in the training data, even in non-swap trials. We also fit BNS to the behavioural output of the trained RNNs to quantify its success in reproducing the target dependence on probe difference (§4). Again, we emphasise that the use of VWM and BNS is one instantiation of our method, and that its ethos of fitting directly to suboptimal behaviour extends beyond this domain, which we discuss briefly in §5.
@@ 3 Methods
Figure 3: DDPM-style trained RNNs accurately captures swap errors in training data, unlike ablated task-optimal networks. A dotted lines are the target swap rate used to generate the synthetic training data for various RNNs, and solid lines are average swap rates inferred from fitting BNS to RNN-generated behaviour after they are trained on this synthetic data (Appendix A; Radmard2025). RNNs achieve fair success in replicating training data for most trials in both probe distance-dependent and -independent cases. Corresponding sample sets in $\mathbb{R}^{2}$ are shown in C, where border colours match the line colours here. B Trajectory snapshots at different points in the second delay period, during which the network is trained to denoise behaviour. Trials are coloured by their argument at the end of denoising, which is interpreted as the colour estimation made by the network. Final samples successfully sample from their multimodal target distribution (Appendix A). C Typical set of final timestep samples for trials with close (left) and far (right) probe feature values. In all cases the red stimulus is cued. Borders indicate generative RNNs, referencing lines in A. Top row (grey border) indicates task-optimal network. D,E Two typical ablations that may be applied to task optimal networks to indcue swap errors: decreasing probe distance beyond the minimum margin between items seen during training to induce confusion, and increasing process noise variance beyond training conditions to increase misselection likelihood. Swap errors would be evidenced by a second mode at the colour of the distractor, indicated by a blue marker. This does not arise in either case, and further attempts would require subjective ablation design. E inset: Conversely, removing process noise for a network trained to perform swap errors recapitulates optimal behaviour.
@@
Our RNN models a population of $n$ densely connected cortical pyramidal neurons [^1] Constants used in defining the architecture/training processes are declared in Appendix G.. Each neuron integrates inputs from a sensory population and all other neurons in the population via a dendritic tree (Appendix B; Lyo2024). Overall, $\mathcal{F}$ in our network dynamics (equation 1) is the somatic integration of the final, most proximal, layer of dendritic nodes with additive Gaussian noise (Appendices B, G). We train the RNN to perform the multiitem delayed estimation task (§2) with 2 items, providing the minimum viable task in which swap errors could arise. The network is first provided with two stimuli, each parameterised by two features, dubbed probe and report (visualised as location and colour), both on a circle. For index-cued networks, only the stimuli colours are provided as individual ordered scalars in $\pi,+\pi)$, and the cue as an index. This provides a direct analogy to the tasks performed by macaques [Panichello2021, Alleman2024], to which we make comparison (§4). Locations are not provided to these networks, and colours are ordered. For feature-cued networks, the stimuli are provided to the network as the conjunctions of locations and colours. These are fed to the network as activations of a palimpsest conjunctively tuned sensory neural population (Appendix B; Matthey2015, Schneegans2017). Importantly, in the latter case, items are order invariant - the network must store the two conjunctions between the two features. This allows us to extend to predictions about the representation of a variable probe feature. The network’s firing rates vector is projected to the 2D plane with a fixed, orthonormal output projection matrix $\bm{x}=\mathcal{G}(\bm{r})=W_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{r}}$$W_{x}\in\mathbb{R}^{2\times n}$ (i.e. $W_{x}^{\intercal}W_{x}=I_{n}$). Network activity is therefore split into two fixed orthogonal subspaces: the behavioural subspace (the network output) $\bm{x}=W_{x}\bm{r}$, and the behavioural nullspace $\bm{m}=W_{x}^{\perp}\bm{r}$. We interpret the argument of this output at the final timestep of the trial, $\arctan\left([W_{x}\bm{r}_{T}]_{1}/[W_{x}\bm{r}_{T}]_{2}\right)$, as the colour recalled by the network on that trial. To train for task performance, a mean squared error (MSE) loss on the discrepancy between this 2D estimate and the report feature of the cued item embedded on the circle in $\mathbb{R}^{2}$ suffices (Appendices A, D).
@@
Training Instead of training for task optimality, we adapt the training procedure used for generative diffusion models (§2; Ho2020, Lyo2024) to train for behavioural realism, accounting for the presence of swap errors. We apply the DDPM training objective only to $\bm{x}$, not the full activity vectofr $\bm{r}$, equating timesteps after cue offset to timesteps in the DDPM denoising process. Besides the use of leaky dynamics (see above), this is a key difference to previous integrations of diffusion models with RNNs in cortical modeling [Lyo2024]. Additionally, the DDPM-style criterion is only applied to the $T$ timesteps following cue offset (Figures 2, 3B) – not repeatedly throughout the full duration of the trial – and the target distribution is trial-specific.

### v2

@@ 1 Introduction
Previous approaches have fallen short of these desiderata. For example, cognitive models – including drift-diffusion [Resulaj et al., 2009, Pardo-Vazquez et al., 2019], Bayesian [Heald et al., 2021], symbolic [Castro et al., 2025], and large language model-based models [Binz et al., 2025] – have been successfully fit to detailed behaviour with impressive predictive power, but their architectures remained too abstract and divorced from RNNs to be useful as hypotheses about neural network dynamics (thus failing Requirement 1). Neural population coding models have suggested causal connections between neural response properties and particular patterns in behaviour [Matthey et al., 2015, Schneegans and Bays, 2017, McMaster et al., 2022], but remained mute about the neural network dynamics that give rise to the hypothesised neural responses in the first place (failing Requirement 1), have been rarely fit to behaviour quantitatively (failing Requirement 2), and their construction was not automated (failing Requirement 3). RNN models – either with hand-crafted [Edin et al., 2009, Bouchacourt and Buschman, 2019], or with task-optimized architectures and parameters [Mante et al., 2013, Stroud et al., 2023] – have been highly successful at suggesting testable hypotheses about neural dynamics underlying cognitive performance, but they rarely capture behaviour beyond generically competent performance, let alone being quantitatively fit to detailed behavioural data (failing Requirements 2, and for hand-crafted networks also Requirement 3). In some cases, non-trivial aspects of behaviour were successfully captured by such models, but at the expense of including purpose-built design choices or ablations [Xie et al., 2023], thus making these models essentially hand-crafted in this sense (failing Requirement 3).
@@
In this work, we introduce an approach that satisfies all three requirements. We extend automated neural mechanism discovery by directly training biologically plausible RNNs to exhibit behavioural suboptimalities seen in experimental data. We tackle two major challenges for doing so. First, RNNs are data hungry, and behavioural data is scarce. To make training feasible, we generate synthetic data using a descriptive generative model, and use this to train our RNN (Figure 1). The second challenge is defining a suitable training objective for the RNN so that they can generate complex (for example, multimodal [Bays et al., 2009] or skewed [Fritsche et al., 2020, Resulaj et al., 2009]) continuous response distributions often found in experiments. Previously used methods for training RNNs are not appropriate for this. Task optimization of RNNs typically used simple optimisation criteria, such as mean squared error or cross entropy [Yang et al., 2019], which encourage unimodal, typically normal response distributions. Even when training RNNs explicitly to generate specific distributions of responses, moment- or score matching-based criteria were used [Echeveste et al., 2020, Chen et al., 2023], which also do not scale well to capturing complex distributions (and require arbitrary choices as to which moments or statistics of the target distribution are important). Instead, by framing elicitation of a behavioural output as generating samples in Euclidean space, we train the RNN using a training criterion inspired by diffusion models [Ho et al., 2020], a state-of-the-art class of generative neural models for complex, continuous distributions. We adapt the training procedure used for diffusion models to work within an RNN, harnessing their flexibility in sampling from complex distributions. This enables us to flexibly fit to continuous response data, unlike prior methods of fitting directly to behaviour [Ji-An et al., 2025], which are limited to categorical distributions (§5).
@@
We demonstrate the power of our approach for automatic mechanism discovery on one of the most studied cognitive paradigms whose neural underpinning is still poorly understood: visual working memory (VWM). While there are many neural circuit models of working memory, they typically do not address the more challenging (and ecologically more relevant) situation when multiple pieces of information (‘items’) need to be maintained in memory Zhang [1996]. Even RNNs that do perform multi-item VWM tasks fail to capture the characteristic patterns of behaviour that are specific to the multi-item situation [Yang et al., 2019, Driscoll et al., 2022], or only capture some select behavioural biases by using hand-crafted networks with purpose-built architectural motifs with limited biological plausibility [Bouchacourt and Buschman, 2019]. In particular, so-called swap errors are a major source of errors that arise when participants recall the wrong item from their working memory, instead of the item that has been cued – and thereby create strongly multimodal response distributions [Bays et al., 2009]. Current RNN models do not capture swap errors, and the resulting multimodal response distributions, at all. Conversely, there exist population coding models that do capture swap errors, but do not make predictions about neural circuit mechanisms; Schneegans and Bays [2017]). We address this gap by showing that our approach produces neural circuits that generate rich, realistic response distributions by directly training networks to perform swap errors. We use our trained networks to generate hypotheses about the biological implementation of observer models from the psychophysics literature (Schneegans and Bays [2016], McMaster et al. [2022]; §4), previously inaccessible to RNNs trained for task performance.
@@ 2 Background
where $\bm{r}$ is the vector of neural firing rates (hidden RNN activations), $\bm{s}$ is the input from an external population (e.g. sensory input into the population), $\bm{\eta}$ is the population’s biological process noise, and $\Lambda$ is the diagonal matrix of neurons’ membrane time constants (typically constant across cells, i.e. $\Lambda=\lambda I$). The non-linear computation $\mathcal{F}$, parameterised by $\theta_{r}$, models the recurrent communication between and information integration within neurons. The network output is generically $\bm{x}=\mathcal{G}(\bm{r};\theta_{o})$, often interpreted as a behavioural response. The exact form of $\mathcal{F}$ and $\mathcal{G}$ vary across studies; in this work, we use a dendritic tree architecture for each neuron (Appendix B; Lyo and Savin [2024]), and interpret the output of the network - a simple linear projection of the firing rates - as a colour estimate on the colour circle. More details are provided in §3, and discussion of our choice of architecture is provided in Appendix B - namely, we chose this architecture to be sufficiently expressive for the task posed, while also maintaining a biologically plausible connectivity structure.
Previous attempts to train RNNs that generate samples typically train RNNs on task-optimality or moment-matching criteria [Echeveste et al., 2020]. This is insufficient for us - as outlined in §1, we are interested in reproducing behaviour that takes on a distinctly multimodal distribution. To do so, we employ methods from the diffusion model literature. We briefly introduce diffusion models here, but defer many details to previous foundational work [Ho et al., 2020].
@@
Neural networks are data-hungry. This is not an obstacle training for task-optimality – task variables such as stimulus features and target responses can be generated procedurally – but it is when training directly on behaviour – producing human or animal behavioural datasets is prohibitively costly at the scales required for training networks. Therefore, we generate synthetic behavioural data to plug this gap. Many of the models listed in §1 which accurately describe behaviour using neural population codes [Schneegans and Bays, 2017], Bayesian cognitive [Heald et al., 2021] and descriptive [Bruijns et al., 2023] models, and/or large language models [Binz et al., 2025] can, by the same token, be used to faithfully generate synthetic behavioural data. These synthetic datasets can be used to train our dynamical neural model; sufficient flexibility in the generative model allows synthetic data that captures desired statistical facets and suboptimalities of behaviour. Many of these models, or their relatives, have also been used to fit behavioural data from working memory (WM) tasks, with varying degrees of satisfaction of our three requirements. For this reason, and because of the sophisticated response distributions elicited from even simple WM tasks, we choose this as the testbed for our method. In §5, we will return to potential extensions to other aspects of WM, and other cognitive tasks in general.
@@
While not consistently distinguished from short-term memory [Aben et al., 2012], WM is generally defined as the temporary maintenance, manipulation, and retrieval of sensory information over an order of seconds, for executive functions. A dominant experimental paradigm in investigating visual working memory (VWM) is the delayed estimation task, in which a human or animal subject is presented with sensory information and asked, after a short delay during which the stimulus is withdrawn, to reproduce some information about it. Figure 2 shows an example of a multiitem delay estimation task, where the subject is presented with multiple (here, 2) items, then presented with a probe dimension (here, location) feature value, called a cue, and asked to reproduce the corresponding report dimension (here, colour) feature value of the cued item. From here, we will interchangeably use ‘location’ and ‘probe feature’, and ‘colour’ and ‘report feature’ for ease of language, however the features used for each role is open for experimental design. Most responses fall near the cued item’s report feature value, and the distribution of estimation errors along the report feature (here, the colour circle) provides key evidence for measuring the capacity of VWM [Ma et al., 2014].
@@
However, there is also a central tendency in responses towards the colour of uncued items, a.k.a. distractors [Bays et al., 2009, Gorgoraptis et al., 2011]. This indicates the presence of swap errors, in which subjects mistakenly recall the colour of a distractor. Behavioural and neural evidence largely implicates misselection of the correct item from memory at cueing time, rather than a misbinding of feature values during memory encoding or storage – swap errors are a failure to extract the correct item from memory, rather than a misencoding into memory (but see Radmard et al. [2025]) Macaque neural signatures reflects this misselection hypothesis Alleman et al. [2024]. The key evidence for this is that swap errors are more likely to happen when a distractor’s probe feature value is close in feature space (here, physical space) to the cued one [Emrich and Ferber, 2012, Schneegans and Bays, 2017, McMaster et al., 2022]. Describing the expected distribution of responses, including swap errors, requires a multimodal distribution, with modes placed over both target and uncued colours.
@@
We use BNS [Radmard et al., 2025] to generate synthetic multiitem delayed estimation data. This model predicts the likelihood of a swap error as a function of the distance of the distractor from the cued stimulus in each feature dimension (Appendix A). In the present work, we do not directly fit BNS to a real behavioural dataset. Instead, we opt to hand-tune the generative model to capture archetypal dependencies of swap errors. Specifically, we train RNNs with synthetic data generated by instantiations of BNS that either i) predicts no swap errors, ii) predicts the same frequency of swap errors for any distractor, and iii) predicts swap errors more frequently when its probe feature is more similar to that of the cued item, as per the extensive experimental evidence. In §3, we show that RNN representations resemble the real macaque neural geometry only in the third case, when maximal realism is captured in the training data, even in non-swap trials. We also fit BNS to the behavioural output of the trained RNNs to quantify its success in reproducing the target dependence on probe difference (§4). Again, we emphasise that the use of VWM and BNS is one instantiation of our method, and that its ethos of fitting directly to suboptimal behaviour extends beyond this domain, which we discuss briefly in §5.
@@ 3 Methods
Figure 3: DDPM-style trained RNNs accurately captures swap errors in training data, unlike ablated task-optimal networks. A dotted lines are the target swap rate used to generate the synthetic training data for various RNNs, and solid lines are average swap rates inferred from fitting BNS to RNN-generated behaviour after they are trained on this synthetic data (Appendix A; Radmard et al. [2025]). RNNs achieve fair success in replicating training data for most trials in both probe distance-dependent and -independent cases. Corresponding sample sets in $\mathbb{R}^{2}$ are shown in C, where border colours match the line colours here. B Trajectory snapshots at different points in the second delay period, during which the network is trained to denoise behaviour. Trials are coloured by their argument at the end of denoising, which is interpreted as the colour estimation made by the network. Final samples successfully sample from their multimodal target distribution (Appendix A). C Typical set of final timestep samples for trials with close (left) and far (right) probe feature values. In all cases the red stimulus is cued. Borders indicate generative RNNs, referencing lines in A. Top row (grey border) indicates task-optimal network. D,E Two typical ablations that may be applied to task optimal networks to indcue swap errors: decreasing probe distance beyond the minimum margin between items seen during training to induce confusion, and increasing process noise variance beyond training conditions to increase misselection likelihood. Swap errors would be evidenced by a second mode at the colour of the distractor, indicated by a blue marker. This does not arise in either case, and further attempts would require subjective ablation design. E inset: Conversely, removing process noise for a network trained to perform swap errors recapitulates optimal behaviour.
@@
Our RNN models a population of $n$ densely connected cortical pyramidal neurons [^1] Constants used in defining the architecture/training processes are declared in Appendix G.. Each neuron integrates inputs from a sensory population and all other neurons in the population via a dendritic tree (Appendix B; Lyo and Savin [2024]). Overall, $\mathcal{F}$ in our network dynamics (equation 1) is the somatic integration of the final, most proximal, layer of dendritic nodes with additive Gaussian noise (Appendices B, G). We train the RNN to perform the multiitem delayed estimation task (§2) with 2 items, providing the minimum viable task in which swap errors could arise. The network is first provided with two stimuli, each parameterised by two features, dubbed probe and report (visualised as location and colour), both on a circle. For index-cued networks, only the stimuli colours are provided as individual ordered scalars in $\pi,+\pi)$, and the cue as an index. This provides a direct analogy to the tasks performed by macaques [Panichello and Buschman, 2021, Alleman et al., 2024], to which we make comparison (§4). Locations are not provided to these networks, and colours are ordered. For feature-cued networks, the stimuli are provided to the network as the conjunctions of locations and colours. These are fed to the network as activations of a palimpsest conjunctively tuned sensory neural population (Appendix B; Matthey et al. [2015], Schneegans and Bays [2017]). Importantly, in the latter case, items are order invariant - the network must store the two conjunctions between the two features. This allows us to extend to predictions about the representation of a variable probe feature. The network’s firing rates vector is projected to the 2D plane with a fixed, orthonormal output projection matrix $\bm{x}=\mathcal{G}(\bm{r})=W_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{r}}$$W_{x}\in\mathbb{R}^{2\times n}$ (i.e. $W_{x}^{\intercal}W_{x}=I_{n}$). Network activity is therefore split into two fixed orthogonal subspaces: the behavioural subspace (the network output) $\bm{x}=W_{x}\bm{r}$, and the behavioural nullspace $\bm{m}=W_{x}^{\perp}\bm{r}$. We interpret the argument of this output at the final timestep of the trial, $\arctan\left([W_{x}\bm{r}_{T}]_{1}/[W_{x}\bm{r}_{T}]_{2}\right)$, as the colour recalled by the network on that trial. To train for task performance, a mean squared error (MSE) loss on the discrepancy between this 2D estimate and the report feature of the cued item embedded on the circle in $\mathbb{R}^{2}$ suffices (Appendices A, D).
@@
Training Instead of training for task optimality, we adapt the training procedure used for generative diffusion models (§2; Ho et al. [2020], Lyo and Savin [2024]) to train for behavioural realism, accounting for the presence of swap errors. We apply the DDPM training objective only to $\bm{x}$, not the full activity vectofr $\bm{r}$, equating timesteps after cue offset to timesteps in the DDPM denoising process. Besides the use of leaky dynamics (see above), this is a key difference to previous integrations of diffusion models with RNNs in cortical modeling [Lyo and Savin, 2024]. Additionally, the DDPM-style criterion is only applied to the $T$ timesteps following cue offset (Figures 2, 3B) – not repeatedly throughout the full duration of the trial – and the target distribution is trial-specific.